# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

### v2

@@ 4.2 Unlocking VLMs with Zero-shot Prompts
We evaluate VLMs in a zero-shot setting, initially focusing on powerful closed-source VLMs (e.g., GPT-4o) to assess whether the GPT4Scene framework can effectively enable 3D scene comprehension. This process is called ”unlocking,” which enables VLMs to understand 3D scenes through prompts without additional training. Specifically, we input $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$. To reduce the cost, we stitch the images from $\bm{V}_{\bm{s}}^{\prime}$ into a single large image. We evaluate three tasks: 3D question answering, dense captioning, and visual grounding. In 3D question answering, the goal is to answer scene-related questions like, ”What is the color of the floor?” In dense captioning, the task is to describe a specific object, such as ”Describe the object represented by $C_{5}$.” In visual grounding, the aim is to identify the object ID from a description, like ”What is the ID of the black chair next to the window?” While question answering is independent of object labels, dense captioning, and visual grounding require object markers. These tasks involve detecting objects and filtering based on the IoU of their bounding boxes. Consistent with Chat-Scene [40] and Robin3D [50], we use Mask3D segmentation results as predicted bounding boxes to calculate the IoU.
In addition to traditional tasks, we have conducted further experiments in this zero-shot setting. Qualitative results are shown in Figure 4. By inputting $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$, VLMs can understand the global features of indoor scenes. At this point, GPT-4o can still accept additional first-person perspective frames, allowing it to understand the current position in the scene to plan the following action. Additionally, using GPT-4o as an agent, VLMs can determine the task type based on the given question and choose the appropriate prompt. Therefore, the GPT4Scene framework shows excellent potential as a core technology for the next generation of embodied intelligence.
@@
Figure 4: Qualitive Results. We demonstrated GPT-4o’s performance with the zero-shot GPT4Scene prompt, enabling it to understand 3D scenes from video frames and BEV images. It can also accept other images and perform various embodied tasks.
@@ 4.3 Enhancing VLMs with ScanAlign Fine-Tuning
Zero-shot prompts can unlock the 3D understanding capabilities of powerful VLMs, but as shown in  Figure 2, this approach does not improve smaller VLMs. Therefore, we aim to enhance open-source, smaller VLMs through fine-tuning. We first construct an indoor scene dataset, ScanAlign, with egocentric, BEV images and text annotations based on the ScanNet [25]. The dataset includes three 3D vision-related tasks represented as $(\bm{V}_{\bm{s}}^{\prime},\bm{I}_{\bm{b}}^{\prime},T)$. The visual input consists of selected video frames with STO markers and BEV images, and $T$ denotes text annotations derived from five ScanNet annotations [25], as shown in Table 1. We use prompts to randomly vary the annotation format to increase annotation diversity, with further details in the supplementary materials. The dataset contains around 165K annotations in total.
@@
After fine-tuning with ScanAlign, during inference, we can input $(\bm{V}_{\bm{s}}^{\prime},\bm{I}_{\bm{b}}^{\prime},Q)$, where $Q$ represents the question. Alternatively, we can inference without the 3D BEV image, using only $(\bm{V}_{\bm{s}}^{\prime},Q)$. For the 3D question answering task, we can even remove all object markers and perform inference using only the original video frames, i.e., $(\bm{V}_{\bm{s}},Q)$, without any additional processing. For the 3D question answering task, we can even remove all object markers, inference with $(\bm{V}_{\bm{s}},Q)$ and use only the original video as input. Our experiments demonstrate that following ScanAlign fine-tuning, small-scale VLMs’ 3D scene understanding capabilities are significantly enhanced.
@@
Vision LLM Based Model  (Close-Source)
GPT-4o [71] 30.9 10.6 15.2 32.6 66.0 40.3 46.1
GPT-4o (GPT4Scene) 35.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+4.7}}}$ $\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.1}}}$ 15.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.4}}}$ 37.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+5.1}}}$ 73.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 42.8$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.5}}}$ 49.4+3.$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+3.3}}}$
Vision LLM Based Model  (Open-Source)
InternVL2-8B [23] 23.9 3.3 14.5 34.3 62.5 33.0 45.3
MiniCPM-V-2.6 [105] 25.1 8.4 11.8 31.5 60.1 42.6 46.6
@@
Qwen2-VL-7B (GPT4Scene)-HDM 44.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.6}}}$ 15.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.5}}}$ 18.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 46.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+17.1}}}$ 96.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+42.4}}}$ 59.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+18.7}}}$ 62.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+15.5}}}$
@@
5 Experiments
In this section, we present the experimental results.  Section 5.1 outlines the implementation details, while  Section 5.2 provides the main results for 3D question answering, dense captioning, visual grounding, and qualitative results. Finally, Section 5.3 covers the ablation study, demonstrating that including global information and STO markers is effective in the training process.
5.1 Implantation Details
All our scene data is based on the ScanNet dataset [25], which contains 1,513 scenes. All these benchmark datasets use the same data split as ScanNet. Our benchmark, as introduced in  Table 1, is divided into three tasks. For the 3D question answering task, we adopt ScanQA [6] and SQA3D [66] as benchmarks. For the 3D dense captioning task, we select Scan2Cap as the benchmark [22]. For the 3D visual grounding task, we use ScanRefer [13] and Multi3DRef [114] to evaluate grounding capabilities. Compared to ScanRefer [13], Multi3DRef evaluates multiple objects, making it more comprehensive and reasonable. We use the same metrics as the original version.
We use the original reconstruction method [26] from ScanNet for reconstruction and align the point cloud. We then use Mask3D [82] to do the 3D instance segmentation, which is consistent with Chat-Scene [40] and Robin3D [50]. After obtaining the 3D segmentation result, we take the xy-coordinates of its center point as the marker’s location in the BEV image. Next, we project the mask onto 2D frames and use the center point of the bounding box of the 2D mask as the marker position. We can directly input the original video during inference to process question answering tasks. However, the markers on 2D images remain essential for dense captioning and visual grounding tasks due to the task’s requirements.
We select $N=8$ frames for each video with a $128\times 123$ resolution. For closed-source VLMs, we use GPT-4o [71], which excels at single images over multiple frames. So, we stitch 8 frames into a large image. For open-source VLMs, the sampled frames are input directly. We use Qwen2-VL-7B-Instruct [90, 117], and in addition to the base setting, we introduce an HD setting (”high resolution”), where the resolution is $512\times 490$. Furthermore, we introduce an HDM setting (”high resolution, more frames”), where the resolution is also $512\times 490$, but with $N=32$ frames. All parameters are fine-tuned using a base learning rate of $5\times 10^{-6}$, cosine annealing, a batch size of 128, and one epoch (1283 steps, 0.1 warmup ratio). Training took approximately six hours on 8 NVIDIA A100 GPUs.
5.2 Main Results
3D Question Answering. The 3D question answering results are shown in Table 2. We categorize the methods into three types: classic task-specific models focused on 3D QA tasks, 3D point LLM-based models, and vision LLM-based models. GPT-4o (GPT4Scene), in a zero-shot manner, outperforms all task-specific models, highlighting the effectiveness of GPT4Scene as the prompt. In contrast, open-source VLMs perform poorly in zero-shot mode without fine-tuning with GPT4Scene, consistent with our pilot study findings. Using the strategy proposed by GPT4Scene, we fine-tune Qwen2-VL-7B, achieving superior question answering results. Notably, Qwen2-VL-7B (GPT4Scene) outperforms all other methods, reaching state-of-the-art performance. Furthermore, compared to the original Qwen2-VL-7B, BLEU-1 in ScanQA improved by 56.1% (27.8 $\xrightarrow{}$ 43.4), and CIDEr increased by 68.6% (53.9 $\xrightarrow{}$ 90.9). In SQA3D, EM-1 rose by 41.0% (40.7 $\xrightarrow{}$ 57.4). Relative to GPT-4o in zero-shot mode, these metrics show gains of 21.9%, 23.7%, and 34.1%, respectively. Our approach greatly improves the model’s understanding of 3D indoor scenes.
3D Dense Caption & Visual Grounding. We also evaluate the model’s performance in dense captioning and visual grounding, which differ from question answering as they require markers to accomplish these tasks. The experimental results are presented in Table 3 and Table 4. As shown, Qwen2-VL-7B, fine-tuned with GPT4Scene, significantly improves VLMs’ 3D captioning and grounding capabilities. Additionally, under high-resolution and larger frame settings (HD), our model achieves state-of-the-art (SOTA) performance, surpassing all existing methods.
@@ 5.2 Main Results
5.3 Ablation Study
In this section, we conduct ablation studies to validate our design modules and analyze factors like resolution and the number of frames.
Ablation on Our Modules. Figure 2 demonstrates that BEV images and STO markers enhance spatial understanding. We use the 3D question answering (QA) task to validate this further because pure video input without STO markers or BEV images can suffice. At the same time, markers are essential for 3D dense caption and visual grounding to reference objects during evaluation. As shown in Table 5, removing BEV images during training and inference reduces performance on both tasks. Further removal of STO markers causes an additional drop in QA performance, highlighting the critical role of BEV images and STO markers in enabling VLMs to understand 3D scenes.
Ablation on Frames and Resolution. Here, we conduct ablation experiments on additional factors using 3D question answering and visual grounding as benchmarks. The results are shown in Table 6. The first three rows show that image resolution significantly affects visual grounding performance but has limited improvement on the QA task. Lastly, the results in the final three rows indicate that increasing the number of frames enhances indoor scene understanding, with a more pronounced effect on grounding performance compared to the limited improvement in QA.
@@ 5.3 Ablation Study
Table 4: Evaluation of 3D Visual Grounding on ScanRefer [13] and Multi3DRef [114]. Our method reaches SOTA performance.
@@
Table 5: Ablation Study. Removing BEV images during training and inference leads to a performance decrease, and further removal of STO markers causes an additional performance decrease.

### v4

@@ 3.2 Unlocking VLMs with Zero-shot Prompts
We evaluate zero-shot VLMs, focusing on powerful closed-source models (e.g., GPT-4o) to test GPT4Scene’s effectiveness in 3D scene understanding. This process is termed “unlocking”, which enables training-free 3D scene comprehension in VLMs via prompts. Specifically, we input $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$. To reduce the cost, we stitch the images from $\mathcal{V}^{*\prime}$ into a single large image. We evaluate three tasks: (1) In 3D question answering, the goal is to answer scene-related questions like, “What is the color of the floor?”. (2) In dense captioning, the task is to describe a specific object, such as “Describe the object represented by $C_{5}$.” (3) In visual grounding, the aim is to identify the object ID from a description, like “What is the ID of the black chair next to the window?” While question answering is independent of object labels, dense captioning, and visual grounding require object markers. These tasks involve detecting objects and filtering based on the IoU of their bounding boxes. Consistent with Chat-Scene [39] and Robin3D [49], we use Mask3D segmentation results as predicted bounding boxes to calculate the IoU.
In addition to traditional tasks, we have conducted further experiments in this zero-shot setting. Qualitative results are shown in Figure 3. By inputting $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$, VLMs can understand the global features of indoor scenes. At this point, GPT-4o can still accept additional first-person perspective frames, allowing it to understand the current position in the scene to plan the following action. Additionally, using GPT-4o as an agent, VLMs can determine the task type based on the given question and choose the appropriate prompt. Therefore, the GPT4Scene framework shows excellent potential as a core technology for the next generation of embodied intelligence.
@@ 3.3 Enhancing VLMs with ScanAlign Fine-Tuning
Zero-shot prompts can unlock the 3D understanding capabilities of powerful VLMs, but as shown in  Table 2, this approach does not improve smaller VLMs. Therefore, we aim to enhance open-source, smaller VLMs through fine-tuning. We first construct an indoor scene dataset, ScanAlign, with egocentric, BEV images and text annotations based on the ScanNet [24]. The dataset includes three 3D vision-related tasks represented as $(\mathcal{V}^{*\prime},\mathcal{I}_{b}^{\prime},T)$. The visual input consists of selected video frames with STO-markers and BEV images, and $T$ denotes text annotations derived from five ScanNet annotations [24], as shown in Table 1. We use prompts to randomly vary the annotation format to increase annotation diversity, with further details in the supplementary materials. The dataset contains 165K annotations in total.
@@
After fine-tuning with ScanAlign, we observed significant performance gains for smaller models, which proves that our approach effectively enhances the intrinsic 3D perception capabilities of visual language models (VLMs). We also conducted a series of ablation experiments to validate this finding.
4 Experiments
Zero-shot 3D QA ROUGE@ScanQA EM-1@SQA3D
VID +GPT4Scene VID +GPT4Scene
Open-sourced VLM Based Model
InternVL2-8B [22] 34.3 33.7-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-0.6}}}$ 33.0 31.4-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-1.6}}}$
MiniCPM-V-2.6-8B [111] 31.5 32.1$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.6}}}$ 42.6 43.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.7}}}$
Qwen2-VL-2B [96] 28.2 28.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.2}}}$ 35.7 34.8-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-0.9}}}$
Qwen2-VL-7B [96] 29.3 31.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.4}}}$ 40.7 41.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+1.0}}}$
[…]
Closed-sourced VLM Based Model
[…]
3D LLM Based Model Pre SOTA Pre SOTA
Chat-Scene [39] 41.6 54.6
Table 2: Zero-shot Setting Results. GPT4Scene yielded limited enhancements for smaller models (2B, 7B, and 8B), while demonstrating significant improvements for larger architectures (72B, closed-source models). This disparity spurred further exploration of fine-tuning strategies for smaller-scale VLMs.
@@ 4 Experiments
Vision LLM Based Model
@@
Qwen2-VL-7B (GPT4Scene)-HDM 44.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.6}}}$ 15.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.5}}}$ 18.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 46.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+17.1}}}$ 96.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+42.4}}}$ 60.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+19.9}}}$ 63.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.4}}}$
@@
In this section, we primarily present the experimental results. First, we evaluate the improvements of GPT4Scene under zero-shot settings in Section 4.1. Following this, we outline the training details in Section 4.2, and subsequently introduce the fine-tuning results in Section 4.3. Finally, Section 4.4 details the ablation study, demonstrating the effectiveness of individual components.
4.1 Zero-shot Setting Results
Can VLMs understand the 3D world directly?
First, we examine GPT4Scene’s zero-shot performance, with the results detailed in Table 2. We begin by comparing a range of open-source models, including InternVL2-8B [22], MiniCPM-V-2.6-8B [111], and the Qwen2-VL family (2B, 7B, and 72B) [96]. Our analysis reveals that for smaller models (2B, 7B, and 8B), the integration of GPT4Scene demonstrates limited effectiveness in enhancing 3D spatial understanding through VLMs. This suggests that the benefits of GPT4Scene are more pronounced in larger models, highlighting the need for further optimization and architectural advancements to bridge this gap.
However, significant improvements are observed with larger models like Qwen2-VL-72B. When testing API-based open-source models Gemini-1.5-Pro [88] and GPT-4o [72], we find that incorporating BEV and STO-markers yields substantial gains, achieving performance comparable to previous state-of-the-art methods Chat-Scene [39]. These results suggest that smaller VLMs face inherent limitations in visual comprehension capabilities, as they cannot directly enhance 3D scene understanding through zero-shot prompting alone. Consequently, we propose constructing a specialized dataset, ScanAlign, and conducting fine-tuning on Qwen2-VL to achieve performance improvements.
4.2 Implementation Details
Our 3D scene understanding benchmark is built upon the ScanNet dataset (1,513 scenes, following its original data splits) and encompasses three core tasks: 3D question answering (evaluated via ScanQA [5] and SQA3D [66]), 3D dense captioning (Scan2Cap [21]), and 3D visual grounding (ScanRefer [12] and Multi3DRef [121]). The data processing pipeline includes reconstructing point clouds with BundleFusion [25] (the original ScanNet method), performing 3D instance segmentation with Mask3D [83] (consistent with Chat-Scene [39] and Robin3D [49]), projecting object centers to bird’s-eye-view coordinates, and mapping 3D masks onto 2D images for marker positioning. 2D marker placement is necessary for dense captioning and visual grounding tasks.
For the experienment config, we sample N=8 frames per video (128×123 resolution). Closed-source visual language models (VLMs) like GPT-4o process stitched 8-frame panoramas, while open-source models (e.g., Qwen2-VL-7B) directly take in sampled frames. The open-source models also support enhanced configurations: HD mode (high resolution, 512×490) and HDM mode (high resolution with multiple frames, 512×490, 32 frames). Training is conducted for one epoch using a base learning rate of 5e-6 with cosine annealing scheduling, completing in approximately 6 hours on 8×A100 GPUs. 2D marker localization is preserved to meet the specific needs of dense captioning and visual grounding tasks.
4.3 Fine-tuning Main Results
We present our experimental results using the ScanAlign dataset for VLM fine-tuning. The results demonstrate that using only Qwen2-VL-7B [96], we achieve state-of-the-art performance across all indoor scene understanding tasks on ScanNet. This validates the effectiveness of GPT4Scene in enhancing VLMs’ capacity for 3D scene comprehension.
3D Question Answering. The 3D question answering results are presented in Table 3. We categorize the evaluated approaches into three groups: (1) classic task-specific models optimized for 3D QA tasks, (2) 3D point cloud LLM-based architectures, and (3) vision-language multimodal LLM systems. Our experimental analysis demonstrates that the baseline Qwen2-VL-7B model without fine-tuning shows constrained capability in 3D QA scenarios. Through systematic fine-tuning with GPT4Scene, the enhanced Qwen2-VL-7B(GPT4Scene) achieves state-of-the-art performance in 3D visual question answering, surpassing all comparative methods across evaluation metrics. Quantitative comparisons reveal 56.1% improvement in BLEU-1 (27.8 → 43.4) and 68.6% enhancement in CIDEr (53.9 → 90.9) on the ScanQA benchmark compared to the original Qwen2-VL-7B implementation. For SQA3D evaluations, we observe a 41.0% increase in EM-1 scores (40.7 → 57.4). Furthermore, under HD (High-Resolution) and HDM (High-Resolution & multi frames) configurations, the model exhibits additional performance gains.
@@ 4.3 Fine-tuning Main Results
Figure 3: Qualitive Results. We demonstrate the qualitative performance of GPT-4o under zero-shot settings using GPT4Scene prompts, highlighting two distinctive application cases that exemplify the versatility of GPT4Scene.
3D Dense Caption & Visual Grounding. We further assess the model’s capabilities in dense captioning and visual grounding tasks, which differ fundamentally from question answering by requiring explicit grounding mechanisms between textual descriptions and 3D spatial structures. As evidenced by the experimental results in Table 4 and Table 6, our GPT4Scene-finetuned Qwen2-VL-7B demonstrates substantial gains in both 3D caption generation and object localization accuracy compared to baseline vision-language models (VLMs). The performance improvements become particularly pronounced under HD and HDM configurations, where our approach establishes new state-of-the-art results across all evaluation metrics, exceeding existing methods by significant margins.
@@
Table 5: Evaluation of 3D Visual Grounding on ScanRefer [12] and Multi3DRef [121]. Our method reaches SOTA performance over all methods.
3D QA GPT Score Combat Situation GPT Score
Methods Win Tie Lose Score
VLM vs 3D LLM
Qwen2-VL-7B [96] vs Chat-Scene [39] 74 243 683 465
Qwen2-VL-7B(GPT4Scene) vs Chat-Scene 543 145 312 1774
Table 6: GPT Score evaluation on 3D question answering. We evaluated the GPT Score on ScanQA [5], comparing our model’s outputs with Chat-Scene [39]. The fine-tuned model achieved significantly higher GPT Scores than current SOTA methods.
GPT Score: New Evaluation Way. We propose a novel GPT Score for 3D QA assessment, comparing two methods with the state-of-the-art 3D LLM Chat-Scene [39] as baseline. In the comparison with Qwen2-VL-7B [96], we input both models’ outputs and the ground truth into GPT-4o [72] to evaluate which response is better. We select 1,000 questions from ScanQA and define ”win,” ”tie,” and ”lose” to indicate whether Qwen2-VL-7B’s answer is superior to, equivalent to, or inferior to Chat-Scene’s [39]. These outcomes are assigned scores of 3, 1, and -1 points respectively (analogous to football scoring), which are aggregated to derive a comprehensive GPT Score. Our results show that Qwen2-VL-7B (without fine-tuning) underperforms Chat-Scene, while its fine-tuned version (GPT4Scene) significantly outperforms it, demonstrating its effectiveness.
Small Objects Sets ScanQA
ROUGE
Chat-Scene 37.5
GPT-4o (GPT4Scene) 35.4
Qwen2-VL-7B (GPT4Scene) 39.4
(a) Ablation on Small Objects.
Precision and Size of STO-markers ScanQA
ROUGE
Qwen2-VL-7B (GPT4Scene) 43.6
- Delete 30% STO-markers 42.7
- Adaptive size adjustment 43.0
(b) Ablation on STO-markers.
BEV Reconstruction Quality ScanQA
ROUGE
Qwen2-VL-7B (GPT4Scene) 43.6
- SLAM3R [63], 50-frame intervals 42.4
- SLAM3R [63], 100-frame intervals 41.9
- SLAM3R [63], 200-frame intervals 43.2
(c) Ablation on BEV Reconstruction Quality.
Table 7: Overall Performance Comparison. We investigated GPT4Scene’s performance on small objects and its robustness in handling STO-markers and BEV reconstruction quality, demonstrating that our model exhibits strong performance and robustness.
@@
Table 8: Ablation Study on our modules. Removing BEV images or STO-markers leads to a performance decrease, and further removal of STO-markers causes an additional decrease.
4.4 Ablation Study
In this section, we conduct ablation studies to validate the effectiveness of GPT4Scene. First, we evaluate its robustness, including performance on small objects, followed by analyzing the robustness of STO-markers and reconstruction quality. Next, we perform module-wise ablation to assess individual components. Finally, we investigate the impact of frame intervals and resolution settings.
Ablation on Robustness Analysis. We tested the performance of GPT4Scene on small objects by selecting 1,000 instances, such as cups and towels. The results demonstrate that the fine-tuned Qwen2-VL-7B [96] model with GPT4Scene exhibits significantly improved performance. Following this, we analyzed the robustness of STO-markers. Two modifications were implemented: first, removing 30% of the markers, and second, dynamically adjusting marker sizes based on object dimensions. Experimental results indicate that both operations had minimal impact on overall performance, confirming the robustness of STO-markers. The third experiment investigated the impact of BEV (Bird’s Eye View) map reconstruction quality on results. Originally, we used BundleFusion [25] from ScanNet [24] for reconstruction. We replaced this with SLAM3R [63], a real-time reconstruction method, and tested settings with frame intervals of 50, 100, and 200. The results show that while reconstruction quality slightly affects performance, the influence is negligible. This suggests that the reconstructed BEV primarily serves to provide global scene context rather than precise geometric details.
Ablation on Our Modules. Table 2 demonstrates that BEV images and STO-markers enhance spatial understanding under zero-shot settings. We further validate this observation in the fine-tuning scenario through evaluations on both 3D question answering and 3D visual grounding tasks. It should be noted that STO-markers are essential for visual grounding——when they are removed, validation can only be performed on the question answering task. As shown in Table 8, both BEV images and STO-markers contribute significantly to the performance improvements.
Ablation on Frames and Resolution. Here, we conduct ablation studies on additional factors using 3D question answering and visual grounding as benchmark tasks. The results are shown in Table 9. We define three configurations: the base mode with 8 frames and longer side resolution of 128, the HD (high-resolution) mode with 8 frames at 512 resolution, and the HDM (high-resolution multi-frame) mode with 32 frames at 512 resolution. The first three rows demonstrate that image resolution substantially impacts visual grounding performance while providing limited improvement for the QA task. The final three rows reveal that frame quantity augmentation enhances indoor scene comprehension, showing more pronounced gains in grounding performance compared to marginal QA improvements.